OpenStudy (anonymous):

how to solve this integral question? integral of [(x power 7m+x power 2m+x power m)].[(2x power 6m+7x power m+14)] power 1/m

OpenStudy (blacksteel):

Is this right? \[\int\limits_{}^{}(x^{7m} + x^{2m} + x^m)*(2x^{6m} + 7x^m + 14)^{1/m}\]

OpenStudy (anonymous):

ya

OpenStudy (blacksteel):

Is the whole equation to the power 1/m, or just the second term?

OpenStudy (anonymous):

only the second term...

OpenStudy (anonymous):

This isn't going to be easy by hand, if at all possible.

myininaya (myininaya):

i suppose we are integrating with respect to x right?

OpenStudy (anonymous):

ya i knw dat...i did dat but amnt getting it..

OpenStudy (blacksteel):

Yeah, I don't think there is an easy way to solve this by hand. Where did the problem come from?

myininaya (myininaya):

this problem is above me lol

myininaya (myininaya):

wait did you try integration by parts?

myininaya (myininaya):

maybe we can find a pattern and right it as a sum

myininaya (myininaya):

write*

OpenStudy (blacksteel):

Integration by parts is useless; Since we don't know what m is, there's no way to reduce the degree of the polynomial.

myininaya (myininaya):

we can see if there is a pattern though and write it as a sum

myininaya (myininaya):

yea i dont think it will work though

OpenStudy (anonymous):

its from tata mc grawls...its realy tough

OpenStudy (anonymous):

all the problem lies in m ... we jus assume that m is one ... that way i can tell what the answer is but with m its too complicated

OpenStudy (anonymous):

@myininaya "we can see if there is a pattern though and write it as a sum" Some fiddling about with Wolfram does appear to give a pattern for the integral: mx^(m-1)[Ax^2m + Bx^6m + Cx^7m + D x^12m + Ex^m + F) all over m(2x^6m + 7x^m + 14) ^ (m-1/m) where A to F are given by A = 35 + 14(m-2) B = 212 +100(m-2) C = 125 +53(m-2) D = 40 14(m-2) E = 77 +35(m-2) F = 28+14(m-2) So you could substitute these in to get the thing but I am feeling too lazy to do that. My guess is Wolfram is using some kind of series calculations so I don't really know how accurate A to F are, maybe more time with the computation would produce something a bit different.

myininaya (myininaya):

yeah i was pretty lazy too about this problem mostly because it looks too ugly

OpenStudy (anonymous):

guys please explain me this question please??

OpenStudy (anonymous):

I thought we did?

OpenStudy (anonymous):

I just put an answer above...

OpenStudy (anonymous):

estudier please explain me step by step please...?

OpenStudy (amistre64):

i recall some place to use a "z sub" where z is an exponential term that plays nicely with others ..

OpenStudy (anonymous):

wats dat?

OpenStudy (amistre64):

cant recall the details a tthe moment. but they take a common factor of the exponents and use a term that can make the work easier by substitution. whether it works here or not i dunno

OpenStudy (anonymous):

wat is z sub?

OpenStudy (amistre64):

its the same as "h sub" but with a z :) like i said, i cant recall the details that well

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I used Wolfram to integrate the expression starting with 1/2 then 1/3, 1/4 etc. From this you can see a pattern in the results from which you can obtain a general result inn m.

OpenStudy (anonymous):

Perhaps there is a tricky way to do it analytically but I don't know it.

OpenStudy (amistre64):

(x^7m^2 +x^2m^2 +1^m) (2x^6m +7x^m +14)^1/m x^14m +x^42m +1^m 2x^6m +7x^m +14 -------------------- 2x^(20m) +2x^(50m) +2x^(7m) 7x^(15m) +7x^(43m) +7x^(2m) +14x^(14m) +14x^(42m) +14x^(m) -------------------------------------------- | 2x^(20m) +2x^(50m) +2x^(7m) | ^1/m | 7x^(15m) +7x^(43m) +7x^(2m) | | +14x^(14m) +14x^(42m) +14x^(m) | ---------------------------------------------- I got know idea what im doing at the moment :)

OpenStudy (anonymous):

u know what happened to Hamilton and the Quaternions, Cantor and his infinities....

OpenStudy (amistre64):

me? aint got clue

OpenStudy (anonymous):

Joke...you will go a bit mad..:-)

OpenStudy (anonymous):

That's not USA mad that's Brit mad.

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

whta u can do is ...take x common in the first part of the question....then take that x into the part which has power .... then substitute the part under power ... u will see the other part gets canceled .....then a simple integration would work

OpenStudy (anonymous):

You lost me....

OpenStudy (anonymous):

u lost me???

OpenStudy (anonymous):

You might be right, I'd like to see it.

OpenStudy (amistre64):

i believe he said factor out an x^m and up it to the mth power and insert it into the mthroot

OpenStudy (amistre64):

i kinda did that by a mth-ed up the whole thing

OpenStudy (anonymous):

no put x into 1/m ..then it becomes x^m

OpenStudy (amistre64):

thats what I said :)

OpenStudy (anonymous):

sry i understood it wrong

OpenStudy (anonymous):

OK, so what's the integral then?

OpenStudy (anonymous):

\[\int\limits_{}^{}(x^{7m-1} + x^{2m-1} + x^{m-1})*(2x^{7m} + 7x^{2m} + 14x^{m})^{1/m}\]

OpenStudy (anonymous):

1/14m{ [(2x^7m + 7x^2m +14x^m)^1/m +1 ] / 1/m + 1}

OpenStudy (anonymous):

\[u=2x^{7m}+7x^{2m}+14x^{m}\] \[1/14\int\limits_{?}^{?}u^{1/m}du\]

OpenStudy (anonymous):

thats what i was trying to say... thank thomas

OpenStudy (anonymous):

You deserve all the credit Ishaan :)

OpenStudy (anonymous):

If you let x = 1/m, what cancels out?

OpenStudy (anonymous):

The way I did it by hand was to first exponentiate everything in the second parentheses by 1/m. Then I expanded everything out. Then break up the integral and do separate u subs.

OpenStudy (anonymous):

you cant let x =i/m ...x is a variable ........while m is const

OpenStudy (anonymous):

its not i its 1

OpenStudy (anonymous):

Oh, i read up a few lines you said to do that?

OpenStudy (anonymous):

no i said to take x common out of the first expression........and then put it under power 1/m

OpenStudy (anonymous):

OH.